On January 1, 2017, McIlroy, Inc., acquired a 60 percent interest in the common stock of Stinson, Inc., for $392,400. Stinson's book value on that date consisted of common stock of $100,000 and retained earnings of $231,900. Also, the acquisition-date fair value of the 40 percent noncontrolling interest was $261,600. The subsidiary held patents (with a 10-year remaining life) that were undervalued within the company's accounting records by $81,700 and an unrecorded customer list (15-year remaining life) assessed at a $57,000 fair value. Any remaining excess acquisition-date fair value was assigned to goodwill. Since acquisition, McIlroy has applied the equity method to its Investment in Stinson account and no goodwill impairment has occurred. At year end, there are no intra-entity payables or receivables.

Intra-entity inventory sales between the two companies have been made as follows:

Year Cost to McIlroy Transfer Price
to Stinson
Ending Balance
(at transfer price)
2017 $130,800 $163,500 $54,500
2018 113,400 151,200 37,800

The individual financial statements for these two companies as of December 31, 2018, and the year then ended follow:

McIlroy, Inc. Stinson, Inc.
Sales $ (741,000 ) $ (377,000 )
Cost of goods sold 487,000 230,200
Operating expenses 199,020 78,400
Equity in earnings in Stinson (35,308 ) 0
Net income $ (90,288 ) $ (68,400 )
Retained earnings, 1/1/18 $ (792,000 ) $ (283,800 )
Net income (90,288 ) (68,400 )
Dividends declared 49,100 19,600
Retained earnings, 12/31/18 $ (833,188 ) $ (332,600 )
Cash and receivables $ 283,600 $ 151,400
Inventory 266,400 132,000
Investment in Stinson 429,006 0
Buildings (net) 347,000 206,500
Equipment (net) 247,700 90,100
Patents (net) 0 24,800
Total assets $ 1,573,706 $ 604,800
Liabilities $ (440,518 ) $ (172,200 )
Common stock (300,000 ) (100,000 )
Retained earnings, 12/31/18 (833,188 ) (332,600 )
Total liabilities and equities $ (1,573,706 ) $ (604,800 )

A. Prepare a consolidated worksheet to determine appropriate balances for external financial reporting as of December 31, 2018.

The debit and credit of consolidation entries, the noncontrolling interest column and the consolidated total column are all necessary in answer.
Answer #1

Part A

Determination of Investment in Stinson account balance

Consideration transferred

392400

Increase in Stinson's retained earnings

31140

Excess fair value amortization

(7182)

2017 ending inventory profit deferral

(10900)

13058

McIlroy's equity earnings in Stinson for 2018

35308

Stinson 2018 dividends declared to McIlroy

(11760)

Investment account balance 12/31/18

$429006

(283800-231900)*60% = 31140

11970*60% = 7182

19600*60% =11760

Acquisition-date fair value allocation and excess amortizations

Consideration transferred

392400

Noncontrolling interest fair value

261600

Subsidiary fair value at acquisition-date

654000

Acquisition-date book value

331900

Fair value in excess of book value

322100

Life in Years

Annual Excess Amortizations

Excess fair value assignments

to patents

81700

10

8170

to customer list

57000

15

3800

to goodwill

183400

indefinite

-

11970

100000+231900 = 331900

322100-81700-57000 = 183400

2017

2018

Intra-entity profits (downstream)

54500

37800

Intra-entity transfers remaining in inventory

20%

25%

Gross profit rate

10900

9450

1-(130800/163500) = 20%

1-(113400/151200)= 25%

Stinson's 2018 income

68400

Excess fair value amortization

(11970)

Adjusted net income

56430

McIlroy's percentage ownership

60%

McIlroy's share of Stinson's adjusted net income

33858

2017 Intra-entity inventory profit recognized

10900

2018 Intra-entity inventory profit deferred

(9450)

McIlroy's equity earnings in Stinson

$35308

Part B

Consolidation Worksheet

Year Ending December 31, 2018

Adjustments & Eliminations

McIlroy

Stinson

Debit

Credit

Noncontrolling interest

Consolidate Totals

Sales

(741000)

(377000)

151200

(966800)

Cost of goods sold

487000

230200

9450

162100

564550

Operating expenses

199020

78400

11970

289390

Equity in earnings of Stinson

(35308)

0

35308

0

Separate company net income

(90288)

(68400)

Consolidated net income

(112860)

To noncontrolling interest

(22572)

22572

To McIlroy, Inc.

(90288)

Retained earnings, 1/1

(792000)

(283800)

283800

(792000)

Net income

(90288)

(68400)

(90288)

Dividends declared

49100

19600

11760

7840

49700

Retained earnings, 12/31

(833188)

(332600)

(833188)

Cash and receivables

283600

151400

435000

Inventory

266400

132000

9450

388950

Investment in Stinson

429006

0

11760

440766

0

Buildings (net)

347000

206500

553500

Equipment (net)

247700

90100

337800

Patents (net)

0

24800

81700

8170

98330

Customer list

57000

3920

53080

Goodwill

183400

183400

Total assets

1573706

604800

2050060

Liabilities

(440518)

(172200)

(612718)

Common stock

(300000)

(100000)

100000

(300000)

Noncontrolling interest 1/1

289422

(289422)

Noncontrolling interest 12/31

(304154)

(304154)

Retained earnings 12/31

(833188)

(332600)

(833188)

Total liabilities and equity

(1573706)

(604800)

925588

925588

(2050060)